# Markov Decision Processes                                  Under External Temporal Processes

### v2

@@ 7.1. Approximate policy evaluation
The best approximate value function $\tilde{V}$ is that function in $\mathcal{F}$ (or equivalently $\mathcal{F}^{T}$) that minimizes the above expected norm difference.
Suppose we are given just $N$ samples $\bar{s}_{1},\dots,\bar{s}_{N}$ from a Markov chain induced by policy $\pi$ in the augmented MDP $\mathcal{M}_{X}$, and corresponding rewards $r_{1},\dots,r_{N}$. The next best thing is to minimize the empirical norm difference at these points, defined as
@@
=$=\left\{\left(f(s^{T}_{1}),\dots,f(s^{T}_{N})\right):f\in\mathcal{F}^{T}\right\}$
@@
where $\Phi=\left(\phi(\bar{s}_{1}^{T}),\dots,\phi(\bar{s}_{N}^{T})\right)_{N\times d}$, and $\bar{s}_{i}^{T}=f^{T}_{trunc}(\bar{s})$.
The true value function $V$ is a fixed point of the Bellman operator $\mathcal{T}$ in $\bar{S}$. From the given $N$ points, we can define our approximate value function as the fixed point of the projected pathwise Bellman operator $\hat{\Pi}\hat{\mathcal{T}}$, where $\hat{T}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is the pathwise Bellman operator, with $\hat{T}(x)$ being defined for $x\in\mathbb{R}^{N}$ as
$\left(\hat{T}(x)\right)_{i}=\begin{cases}r_{i}+\gamma x_{i+1}&\text{ for }i<N\\ r_{i}&\text{ for }i=N\end{cases},$
@@
$\hat{\Pi}(x)=\mathop{\mathrm{argmin}}\limits_{z\in\mathcal{F}_{N}}\left|\left|x-z\right|\right|_{N}^{2}.$
This also be written as $\hat{T}=r+\gamma\hat{P}$, where $r=(r_{t})_{t\in[N]}$ and $\hat{P}$ is the path-wise transition operator. $\hat{T}$ is a $\gamma$-contraction in ||.$||.||_{N}$ as well as ||.$||.||_{2}$, and the projection operator $\hat{\Pi}$ onto $\mathcal{F}_{N}$ is non-expansive since $\mathcal{F}$ is convex.
This means that $\hat{\Pi}\hat{T}$ is a $\gamma$-contraction on $\mathbb{R}^{N}$ and has a unique fixed point $\hat{v}\in\mathcal{F}_{N}$, which we define as our approximation of the value function values at these points. Corresponding to this point, there exists function $\hat{V}\in\mathcal{F}$ such that $\hat{v}=\left(\hat{V}(\hat{s}_{t})\right)_{t\in[N]}$.
@@
This $\hat{V}$ is considered the output of the approximate policy evaluation step of the policy iteration algorithm. For determining the overall error of this estimate, the intermediate task is to bound the difference between this and the true solution on these data points, i.e, $||V-\hat{V}||_{N}$, as a function of the number of data points and the complexity of the function class under consideration.
@@
$||V-\hat{V}||_{N}=||v-\hat{v}||_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{1}{1-\gamma}||\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v||_{N}.$
The first term is the approximation error due to having a restricted function space and is unavoidable. The second term also includes the estimation error due to using the pathwise Bellman operator instead of the actual Bellman operator, and is a function of the number of samples used. It can be written as $||\hat{\Pi}\xi||_{N}$, where $\xi=\hat{\mathcal{T}}v-v$ is the estimation error. Now, for $t<N$,
@@
$\xi_{t}$ are functions of $\bar{s}_{t}$, that are samples from the Markov chain induced by the policy $\pi$ in the augmented MDP $\bar{M}$. Considering them as a random variables, $\left(\xi_{t}\varphi_{t}(\bar{s_{t}})\right)_{1\leq t<N}$ is a Martingale difference sequence w.r.t $\bar{s}_{t}$, with each element being zero mean and bounded by $\frac{L}{1-\gamma}$. So, applying Azuma’s inequality to $(\xi_{t}\varphi_{i}(\bar{s})_{t})_{1\leq t<N}$ gives, for each $i\in[d]$,
@@
Following Lazaric et al. (2012), this can be used to bound $||\hat{\Pi}\xi_{n}||_{N}$ as
$\left|\left|\hat{\Pi}\hat{\mathcal{T}}v-\hat{\Pi}v\right|\right|_{N}\leq\frac{1}{N}\sqrt{\frac{d}{\nu_{n}}}\max_{i}\left|\sum_{t=1}^{N}\xi_{t}\varphi_{i}(s_{t})\right|,$
@@
$||v-\hat{v}||_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{1}{1-\gamma}||\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v||_{N}$
$\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu_{N}}}\left(\sqrt{\frac{2\log(2d/\delta)}{N}}+\frac{1}{N}\right).$ (2)
The above identity is almost identical to the one in Lazaric et al. (2012), the reason being that we are still operating in the range space of the basis functions restricted to a finite set of states and so the structure of the MDP in our setting does not affect the analysis. Since the proof follows a similar path, we have only included those steps where there is a discrepancy, which are mainly due to the noise terms taken in the martingale difference sequence. In our setting, these noise terms also include stochasticity due to the reward in addition to the transition function, leading to a difference in the final expression.
While the final inequality is quite similar, we shall see that when considering the expected error in the augmented state space, we can bring out the properties of our setting by splitting the expected error in terms of error due to function truncation and the inherent error due to linear function approximation.
@@ 7.2. Expected error
Let $\bar{s}_{1},\dots,\bar{s}_{N+\tilde{N}}$ be a sample path obtained using policy $\pi$ in the MDP $\mathcal{M}_{X}$. Suppose sufficiently large $\tilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ are used to compute the truncated least-squares path-wise estimate $\tilde{V}$ of the true value function $V$. Let $\nu$ be a lower bound on the eigenvalues of the sample-based Gram matrix that hold with probability $41-\delta/$ 4. Then, with probability at least $1-\delta$,
$||\tilde{V}-V||_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(8d/\delta)}{N}}+\frac{1}{N}\right)+\bar{\epsilon}(N,d,\delta).$
Proof.
Following proof of Theorem $5$ of Lazaric et al. (2012),
$2||\hat{V}-V||_{N}\geq 2||\tilde{V}-V||_{N}\geq||\tilde{V}-V||_{\rho}-\epsilon_{1},$
where the first inequality is because of truncation, and the second inequality is a generalization Lemma for Markov chains, and so
$||\tilde{V}-V||_{\rho}$ $\leq 2||\hat{V}-V||_{N}+\epsilon_{1}$
$\leq\frac{2}{\sqrt{1-\gamma^{2}}}||V-\hat{\Pi}V||_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
$\leq\frac{2}{\sqrt{1-\gamma^{2}}}||V-\Pi V||_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}||V-\Pi V||_{\rho}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-\Pi^{trunc}V||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
@@
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-V^{\epsilon}(s,x_{0:T},\bar{0})||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
@@
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-V^{\epsilon}||_{\rho}+||V^{\epsilon}-V^{\epsilon}(s,x_{0:T},\bar{0})||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
@@
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
@@
The Lemmas used in the proof can be taken from Lazaric et al. (2012) without any modifications because, even if some of the results, specifically the results on the bound on the covering numbers for linear function spaces taken from Györfi et al. (2002), require the domain of the basis functions to be finite dimensional, the analysis takes place in the co-domain and so can be directly applied to our setting. We state the complete theorem along with the Lemmas in Appendix B.
@@ 8. Experiments
We conducted experiments for solving a non-stationary variant of classic Pendulum-v1 environment in the control suite of OpenAI Gym. The task consists of applying torque to a pendulum to swing it and keep it upright. We considered a fixed neural-network policy that has been pre-trained partially using DDPG. We then used pathwise LSTD with linear function approximation for evaluating this policy as described in Section 7. The features considered are the standard cosine and sine of the angle and the angular velocity, with additional features being the angle itself along with the squares of all these features, making the state $8$ dimensional.
@@
For a fixed event horizon $5$, the error reduces as the number of samples increases, which is quite intuitive. For a fixed number of samples $10000$, the average error initially reduces as the event horizon increases, corresponding to the first term in the sample complexity of the total variation induced by events older than the event horizon. After a certain stage, increasing the event horizon results in a gradual increase in the error. This is because the events older than the horizon now contribute a negligible influence on the current dynamics, while the dimensionality of the features is increasing due to the inclusion of more past events in the features, resulting in a slightly greater expected error. The third plot shows the dependence on $\lambda_{\alpha}$, the rate of decay of $\alpha_{t}$. Faster decay results in lesser non-stationarity and lower approximation error due to the first term in the sample complexity upper bound. Due to high variation in the results obtained, we conducted $20$ trials and plotted the median of the Mean Squared Errors for the experiments w.r.t decay rate and number of samples. The error bars indicate the range of $40$-$60$ percentile versus the decay rate and $20$-$80$ percentile for the other two.

### v3

@@ 6.1. Approximate policy evaluation
The best approximate value function $\widetilde{V}$ is that function in $\mathcal{F}$ (or equivalently $\mathcal{F}^{(T)}$) that minimizes the above expected norm difference.
Suppose we are given just $N$ samples $\bar{s}_{1},\dots,\bar{s}_{N}$ from a Markov chain induced by policy $\pi$ in the augmented MDP $\mathcal{M}_{X}$, and corresponding rewards $r_{1},\dots,r_{N}$. We minimize the empirical norm difference at these points, defined as
@@
=$=\left\{\left(f(s^{T}_{1}),\dots,f(s^{T}_{N})\right):f\in\mathcal{F}^{(T)}\right\}$
@@
where $\Phi=\left(\phi(\bar{s}_{1}^{T}),\dots,\phi(\bar{s}_{N}^{T})\right)_{N\times d}$ and $\bar{s}_{i}^{T}=f^{(T)}_{trunc}(\bar{s})$.
The true value function $V$ is a fixed point of the Bellman operator $\mathcal{T}$ in $\bar{S}$. From the given $N$ points, we can define our approximate value function as the fixed point of the projected pathwise Bellman operator $\hat{\Pi}\hat{\mathcal{T}}$, where $\hat{\mathcal{T}}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is the pathwise Bellman operator, with $\hat{\mathcal{T}}(x)$ being defined for $x\in\mathbb{R}^{N}$ as
$\left(\hat{\mathcal{T}}(x)\right)_{i}=\begin{cases}r_{i}+\gamma x_{i+1}&\text{for }i<N\\ r_{i}&\text{ for }i=N\end{cases}$
@@
$\hat{\Pi}(x)=\mathop{\mathrm{argmin}}\limits_{z\in\mathcal{F}_{N}}\left\|x-z\right\|_{N}^{2}.$
This also be written as $\hat{\mathcal{T}}=r+\gamma\hat{P}$, where $r=(r_{t})_{t\in[N]}$ and $\hat{P}$ is the path-wise transition operator. $\hat{\mathcal{T}}$ is a $\gamma$-contraction in .$\|.\|_{N}$ as well as .$\|.\|_{2}$, and the projection operator $\hat{\Pi}$ onto $\mathcal{F}_{N}$ is non-expansive since $\mathcal{F}$ is convex.
This means that $\hat{\Pi}\hat{\mathcal{T}}$ is a $\gamma$-contraction on $\mathbb{R}^{N}$ and has a unique fixed point $\hat{v}\in\mathcal{F}_{N}$, which we define as our approximation of the value function values at these points. Corresponding to this point, there exists function $\hat{V}\in\mathcal{F}$ such that $\hat{v}=\left(\hat{V}(\hat{s}_{t})\right)_{t\in[N]}$.
@@
This $\hat{V}$ is considered the output of the approximate policy evaluation step of the policy iteration algorithm. For determining the overall error of this estimate, the intermediate task is to bound the difference between this and the true solution on these data points, i.e, $\|V-\hat{V}\|_{N}$, as a function of the number of data points and the complexity of the function class under consideration.
@@
$\left\|V-\hat{V}\right\|_{N}=\left\|v-\hat{v}\right\|_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}\left\|v-\hat{\Pi}v\right\|_{N}+\frac{1}{1-\gamma}\left\|\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v\right\|_{N}.$
The first term is the approximation error due to having a restricted function space and is unavoidable. The second term also includes the estimation error due to using the pathwise Bellman operator instead of the actual Bellman operator, and is a function of the number of samples used. It can be written as $\left\|\hat{\Pi}\xi\right\|_{N}$, where $\xi=\hat{\mathcal{T}}v-v$ is the estimation error. Now, for $t<N$,
@@
$\xi_{t}$ are functions of $\bar{s}_{t}$, that are samples from the Markov chain induced by the policy $\pi$ in the augmented MDP $\mathcal{M}_{X}$. Considering them as a random variables, $\left(\xi_{t}\varphi_{t}(\bar{s_{t}})\right)_{1\leq t<N}$ is a Martingale difference sequence w.r.t $\bar{s}_{t}$, with each element being zero mean and bounded by $\frac{L}{1-\gamma}$. So, applying Azuma’s inequality to $(\xi_{t}\varphi_{i}(\bar{s}_{t}))_{1\leq t<N}$ gives, for each $i\in[d]$,
@@
Following Lazaric et al. (2012), this can be used to bound $\|\hat{\Pi}\xi_{n}\|_{N}$ as
$\left\|\hat{\Pi}\hat{\mathcal{T}}v-\hat{\Pi}v\right\|_{N}\leq\frac{1}{N}\sqrt{\frac{d}{\nu_{n}}}\max_{i}\left|\sum_{t=1}^{N}\xi_{t}\varphi_{i}(s_{t})\right|,$
@@
$\left\|v-\hat{v}\right\|_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}\left\|v-\hat{\Pi}v\right\|_{N}+\frac{1}{1-\gamma}\left\|\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v\right\|_{N}$
$\leq\frac{1}{\sqrt{1-\gamma^{2}}}\left\|v-\hat{\Pi}v\right\|_{N}+\frac{L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu_{N}}}\left(\sqrt{\frac{2\log(2d/\delta)}{N}}+\frac{1}{N}\right).$ (1)
The above identity is almost identical to the one in Lazaric et al. (2012), the reason being that we are still operating in the range space of the basis functions restricted to a finite set of states and so the structure of the MDP in our setting does not affect the analysis. Since the proof follows a similar path, we have only included those steps that differ from the original proof, which are mainly due to the noise terms taken in the martingale difference sequence. In our setting, these noise terms also include stochasticity due to the reward in addition to the transition function, leading to a difference in the final expression.
While this final inequality is quite similar, we shall see that when considering the expected error in the augmented state space, we can bring out the properties of our setting by splitting the expected error in terms of error due to function truncation and the inherent error due to linear function approximation.
@@ 6.2. Expected error
Let $\bar{s}_{1},\dots,\bar{s}_{N+\widetilde{N}}$ be a sample path obtained using policy $\pi$ in the MDP $\mathcal{M}_{X}$. Suppose sufficiently large $\widetilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ samples are used to compute the truncated least-squares path-wise estimate $\widetilde{V}$ of the true value function $V$. Let $\nu$ be a lower bound on the eigenvalues of the sample-based Gram matrix that holds with probability $41-\delta/$ 4. Then, with probability at least $1-\delta$,
$\left\|\widetilde{V}-V\right\|_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(8d/\delta)}{N}}+\frac{1}{N}\right)+\bar{\epsilon}(N,d,\delta),$
where $V$ is the true value function, $\widetilde{V}$ is the learned value function clipped at $\frac{1}{1-\gamma}$, $\Pi^{trunc}V$ is the best possible value function on the truncated state space, $\Pi V$ is the best approximating value function in the linear function space considered, $M_{t}$ and $N_{t}$ are the upper bounds on the total variation due to exogenous events older than $t$ time steps induced in the transition dynamics and event mark distribution respectively, and $\bar{\epsilon}$ depends on the mixing rate of the Markov chain, dimensionality $d$ of the features, and the probability of failure $\delta$.
This result breaks down the overall error into individual components, viz. the approximation error due to the non-stationarity, the inherent error due to linear function approximation, the error due to stochasticity of the samples that reduces with the number of samples, and other errors due to mixing of the Markov chain, the dimensionality of the function space, etc. The detailed conditions on the MDP under which this result holds, as well as the exact expression for $\bar{\epsilon}$, are given in the Appendix.
Proof of Theorem 6.1.
Following proof of Theorem $5$ in Lazaric et al. (2012),
$2\left\|\hat{V}-V\right\|_{N}\geq 2\left\|\widetilde{V}-V\right\|_{N}\geq\left\|\widetilde{V}-V\right\|_{\rho}-\epsilon_{1},$
where the first inequality results from truncation, and the second inequality is a generalization Lemma for Markov chains, and so
$\left\|\widetilde{V}-V\right\|_{\rho}$ $\leq 2\left\|\hat{V}-V\right\|_{N}+\epsilon_{1}$
$\leq\frac{2}{\sqrt{1-\gamma^{2}}}\left\|V-\hat{\Pi}V\right\|_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
$\leq\frac{2}{\sqrt{1-\gamma^{2}}}\left\|V-\Pi V\right\|_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left\|V-\Pi V\right\|_{\rho}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\left\|V-\Pi^{trunc}V\right\|_{\rho}+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
@@
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\left\|V-V^{\epsilon}(s,x_{0:T},\bar{0})\right\|_{\rho}+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
@@
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\left\|V-V^{\epsilon}\right\|_{\rho}+\left\|V^{\epsilon}-V^{\epsilon}(s,x_{0:T},\bar{0})\right\|_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
@@
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
@@
The Lemmas used in the proof can be taken from Lazaric et al. (2012) without any modifications because, even if some of the results, specifically the results on the bound on the covering numbers for linear function spaces taken from Györfi et al. (2002), require the domain of the basis functions to be finite dimensional, the analysis takes place in the co-domain and so can be directly applied to our setting. We state the complete theorem along with the Lemmas in Appendix C.
@@ 7. Experiments
We conducted experiments for analyzing policy evaluation in a non-stationary variant of the classic Pendulum-v1 environment in the control suite of OpenAI Gym. The task consists of applying torque to a pendulum to swing it and keep it upright. We considered a fixed neural-network policy that has been pre-trained partially using DDPG (Lillicrap et al., 2016). We then used pathwise LSTD with linear function approximation for evaluating this policy as described in Section 6. The features considered are the standard cosine and sine of the angle and the angular velocity, with additional features being the angle itself along with the squares of all these features, making the state $8$ dimensional.
@@
For a fixed event horizon $5$, the error reduces as the number of samples increases, which is quite intuitive. For a fixed number of samples $10000$, the average error initially reduces as the event horizon increases, corresponding to the first term in the sample complexity of the total variation induced by events older than the event horizon. After a certain stage, increasing the event horizon results in a gradual increase in the error. This is because the events older than the horizon now contribute a negligible influence on the current dynamics, while the dimensionality of the features is increasing due to the inclusion of more past events in the features, resulting in a slightly greater expected error.
The first plot shows the dependence on $\lambda_{\alpha}$, the rate of decay of $\alpha_{t}$. Faster decay results in lesser non-stationarity and lower approximation error due to the first term in the sample complexity upper bound.
Due to high variation in the results obtained, we conducted $20$ trials and plotted the median of the Mean Squared Errors for the experiments w.r.t decay rate and number of samples. The error bars indicate the range of $40$-$60$ percentile versus the decay rate and $20$-$80$ percentile for the other two.
These results empirically demonstrate the sample complexity bound for policy evaluation with linear function approximation in Theorem 6.1. The average error decreases with increasing $N$ and decreasing $N_{t}$.
Increasing the event horizon $T$ increases $d$ while reducing $\sum_{t=T+1}^{\infty}\{M_{t},N_{t}\}$, resulting in a trade-off between the corresponding two terms in the bound, with the minimal error being achieved for $T=2$.
8. Related Work
Many previous studies in the literature that address non-stationarity primarily provide practical solutions that lack theoretical backing. While some research does offer theoretical insights in specific special cases, our work tackles this problem from a comprehensive perspective. The nature of the results reported in this paper is significantly different from those available in the existing literature.
8.1. Practical approaches in the literature
Numerous previous studies examine environments characterized by piece-wise stationarity, where arbitrary changes transpire at designated change points. Different studies adopt differing assumptions regarding the prior knowledge accessible to the agent concerning the nature of these change points. For instance, Padakandla et al. (2020) assume that the transition and reward functions belong to a fixed set of ‘contexts’ between change points. This work introduces a model-free algorithm that employs a change point detection mechanism to ascertain the timing changes by analyzing the state-reward-next state tuples obtained from the environment. Although this can be regarded as a special case within our framework, it is important to note that the context remains unobserved in this scenario, while the pattern of change in the context is presumed to be known a priori. Further, they consider only a fixed set of potential contexts. Due to these issues, the proposed algorithm in Padakandla et al. (2020) cannot be applied in our case since we consider a continuous space of potential contexts and no knowledge of the order in which the contexts will occur.
Alegre et al. (2021) proposed a model-based algorithm that can deal with arbitrarily different types of contexts and continuous state and action spaces. Although this approach considers a relatively more general setting, the focus of this work is to detect the change points efficiently and to learn separate models for each context. This is infeasible for our setting where the context changes possibly every time step.
Hadoux et al. (2014) formalized a class of piece-wise stationary MDPs called Hidden-Semi-Markov-Model MDPs where the transition from one hidden stationary MDP mode to another is semi-Markov in nature. To solve this problem, this work adapts the Partially Observable Monte Carlo Planning (POMCP) algorithm used for solving Partially Observable Markov Decision Processes (POMDPs). This work, too, does not provide any theoretical results.
8.2. Theoretical results available in the literature
Lecarpentier and Rachelson (2019) considers a finite horizon MDP in which the transition dynamics and reward function evolve such that they are Lipschitz continuous with respect to time. Cheung et al. (2020) considers such a setting in a finite MDP where the amount of total change in the environment, referred to as the variation budget, is known to the agent. This work proposes an optimistic Value Iteration based algorithm and provides guarantees on the upper bound of its dynamic regret.
Wei and Luo (2021) proposes a black-box reduction that turns any suitable algorithm with optimal performance in a near-stationary environment into an algorithm that achieves optimal dynamic regret in a non-stationary environment.
Some algorithms assume additional structure on the nature of non-stationarity. Feng et al. (2022) propose Factored Non-stationary MDPs (FN-MDPs), where the transition dynamics are defined in terms of a Dynamic Bayesian Network over the states, actions, rewards, and latent change factors that cause non-stationarity. Here, both the states and the latent change factors have transition dynamics obeying this causal graph, and the latent factors evolve in a Markov fashion. They propose using a Variational Auto-Encoder (VAE) framework to learn transition dynamics and present results in MuJoCo and robotic manipulation tasks.
A factorization setting with a resemblance to ours is the Exogenous MDP (Efroni et al., 2022), wherein the environment state is divided into two parts, an endogenous state which is controllable by the agent’s actions, and an exogenous state, which evolves on its own and does not affect the agent’s reward. The precise way in which this division exists is unknown to the agent. While this may seem similar to our formulation of states and external events, in our case, the external events do influence the dynamics of the state and, thereby, the rewards obtained by the agent. Further, Efroni et al. (2022) only consider the tabular setting and propose algorithms that can only deal with finite horizon episodes.
A similar decomposition of the state space into endogenous and exogenous sub-spaces was studied in Dietterich et al. (2018), where it is assumed that the exogenous states evolve in a Markov fashion and the reward function can be additively decomposed into an exogenous reward and an endogenous reward. This work proposed algorithms to determine the exogenous component of the state as projections that maximize the partial correlation coefficient as a proxy for the mutual information between the exogenous next state and the endogenous current state and action conditioned on the exogenous current state.
8.3. Positioning our work in the literature
While previous studies like Efroni et al. (2022); Dietterich et al. (2018) focus on the decomposition of states into endogenous and exogenous components, our work addresses a fundamentally different problem: the impact of integrating exogenous event information into reinforcement learning.
Changing dynamics of environments based on externally specified “contexts” has also been studied under Contextual MDPs (Hallak et al., 2015). However, these studies consider contexts, where each context is constant for each episode and has a finite number of possible values.
The aforementioned work proposes the CECE algorithm that learns a set of policies corresponding to all the contexts and determines the latent context for an episode (if unknown) to choose the policy corresponding to the current context.
More specifically, CECE clusters an initial set of trajectories into groups corresponding to a fixed set of latent contexts. For each new episode, a partial realization is used to classify it into one of these clusters and to use a model learned on that cluster to exploit the rest of the trajectory. This algorithm cannot handle the setting considered in this paper because each new time step results in a change in the context, and knowledge of a latent context for the current time step does not guarantee the same context for the rest of the episode, except in a trivial special case.
Modi et al. (2018) consider a relatively closer setting to ours where the context is known and possibly continuous, with their algorithm being more amenable to being adapted to handle dynamic contexts. However, it works by keeping count of the number of occurrences of $(s,a)$ in each of a set of representative contexts to learn a set of explicit models, which is only possible in the setting of finite state and action spaces.
Tennenholtz et al. (2023) extend Contextual MDPs to handle dynamic contexts that are known to the agent, change at each time step, and influence the state transition distribution. However, this work differs from ours in several significant aspects. This work considers the contexts at each time step to depend on the history of states, actions, and contexts, while the next state depends only on the current state, action, and context. This means causality exists in both directions, from states to contexts and vice versa. Whereas in our work, the contexts are exogenous events that cannot be controlled by the agent, and the causality flows only in one direction, from contexts to states.
Further, this work considers finite state spaces, and their analysis is explicitly geared towards such spaces. Their bounds depend on the size of the state and action sets.
In Tennenholtz et al. (2023), the context space is a fixed finite set of $M$ vectors, and a latent map is learned from the (state, action, context) space to $\mathbb{R}^{M}$, with the range of the map still being a finite albeit exponentially increasing set. This work also considers a finite horizon MDP. This relatively simple setting allows one to tractably find the latent map and learn an optimal policy. In this sense, our setting is much more general and challenging. In addition to this, the work of Tennenholtz et al. (2023) presents regret analysis, while our study focuses on the convergence results of algorithms.
Our analysis addresses the non-stationarity induced in the MDP due to the influence of non-Markovian events, necessitating the learning of policies that are contingent upon both the current state and a history of prior events. Although this approach bears resemblance to the strategies employed in Partially Observable MDPs (POMDPs) (Sunberg and Kochenderfer, 2018), the underlying rationale diverges significantly.
In POMDPs, histories of observations are utilized despite the Markovian nature of state transition and observation functions, primarily to mitigate the agent’s inability to perceive the true state. Conversely, in our framework, both the state and external events are fully observable, and the necessity for an augmented space that incorporates event histories arises from the intrinsic non-Markovian characteristics of the non-stationary environment.
Moreover, the causal framework of our scenario is distinctly different. The events exhibit non-Markovian properties, and the state at any given time directly affects the distribution of the subsequent state, regardless of the knowledge of the event. In POMDPs, while beliefs and sufficient statistics are contingent upon a history of observations, the states themselves maintain a Markovian nature, leading to observations that are causally independent.
Our paper makes significant inroads into understanding how characteristics of exogenous event processes affect the tractability of the problem and the convergence and sample complexity of RL algorithms. To the best of our knowledge, there are no existing works in the literature that offer these findings.